OpenStudy (anonymous):

Can anybody explain how to get the answer for: A book is being pushed along a table at constant speed. The coefficient of static friction is 0.75, and the coefficient of kinetic friction is 0.30. The book has a mass of 0.50 kg. What is the force of friction?

OpenStudy (ybarrap):

 Since the book is "moving" we only need to use the kinetic friction coefficient. All forces must sum to zero. Let \(F\) be the force applied to push the book, and \(f_f\) be the frictional force of the table against the book. $$ F-f_f=0\\ F=f_f\\ $$ The kinetic frictional force \(\mu_k=0.3\) is a fraction of the weight of the book: $$ F=\mu_k\text{mass}\times g $$ Where \(g\) is the acceleration of gravity \(9.1~m/s^2\). So, $$ F=0.3\times0.5\times9.81~N $$ Since \(f_f=F\), this will also be the force of friction. Make sense?